BP: Given the following differential equation, solve and design the RLC Circuit. Y + 5y + 6y = 72
Answer #1

first we solve the differential equation using laplace transform and then comparing it with series RLC circuit transfer function we get value of R , L and C .

ÿ + sý+by = 72 using laplace transform on the 2nd order differential ef & considering tero initial Cand - šY (31 +5 sy(3) + 6DULU2 PIVE Vo(s)= df (3) R + Lst - RestLešti and 72 US 12 .6 5+55+6 st 55+6 = comparing ofw - Te = 6 taking, R = b - L = c =so, R-L-c circuit mm σου σο I leuren en = 1666.6 MF (V.) 12 uit Ult) son 107 MEES